# Flow-Anchored Consistency Models

### v1

@@ 5 Limitations and Future Work
Our primary limitation is the performance gap between 2-step and 1-step generation, indicating room to improve 1-step expressiveness. We also observe a trade-off in multi-step performance: according to Figure 3(c), after reaching the optimal 2-step result, further training continues to improve the 1-step FID, while slightly degrading the 2-step result. We hypothesize this is due to a slight deviation between the intermediate samples generated by our multi-step sampler and the true flow trajectory. Exploring sampling strategies that better align with the true flow is a promising direction for future work. Additionally, the distillation stage is computationally intensive. Each training step requires two forward passes of the teacher model for CFG, a standard forward pass of the student model, and a separate Jacobian-vector product (JVP) forward pass. This JVP computation constitutes a significant bottleneck as it is not readily accelerated by modern compilers like torch.compile. Finally, we plan to further validate our approach on larger-scale models and more complex conditional generation tasks, such as text-to-image synthesis.
@@ 6 Conclusion
We identify that the training instability of continuous-time Consistency Models stems from a “missing anchor” in the underlying velocity field. We introduce the Flow-Anchored Consistency Model (FACM), a simple, architecturally-agnostic training strategy that solves this problem by using a Flow Matching loss as an explicit anchor, and present two effective implementations for this mixed-objective design. Our method achieves new state-of-the-art FIDs on both ImageNet 256$\times$256 (1.76 at NFE=1 and 1.32 at NFE=2) and CIFAR-10 (2.69 at NFE=1 and 1.87 at NFE=2). FACM provides a principled and effective recipe for building high-performance, few-step generative models.
@@ Appendix A Appendix
A.1 On Total Derivatives
@@
In this paper, for a network $\bm{F}_{\theta}(\bm{x}_{t},\mathcal{C}(t))$, its total derivative along the trajectory $\bm{x}_{t}(t)=(1-t)\bm{x}_{0}+t\bm{x}_{1}$ with respect to $t$ is given by the chain rule:
$\frac{\mathrm{d}\bm{F}_{\theta}(\bm{x}_{t},\mathcal{C}(t))}{\mathrm{d}t}=\frac{\partial\bm{F}_{\theta}}{\partial\bm{x}_{t}}\bm{v}+\nabla_{\mathcal{C}}\bm{F}_{\theta}\cdot\frac{\mathrm{d}\mathcal{C}(t)}{\mathrm{d}t}.$ (16)
@@
The term $\frac{\mathrm{d}\bm{F}_{\theta}(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}$ is computed for the CM task. Depending on the implementation strategy (Sec. 3.2.2), the conditioning $c_{\text{CM}}$ can be $t$ or a tuple $(t,1)$. In both cases, its derivative with respect to $t$ is effectively 1 for the time-dependent component and 0 for any constant component. Therefore, the calculation simplifies to:
$\frac{\mathrm{d}\bm{F}_{\theta}(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}\approx\frac{\partial\bm{F}_{\theta}}{\partial\bm{x}_{t}}\bm{v}+\frac{\partial\bm{F}_{\theta}}{\partial t},$ (17)
@@
where $\frac{\partial\bm{F}_{\theta}}{\partial t}$ denotes the partial derivative with respect to the explicit time argument(s) encoded in the conditioning. This term can be calculated via JVP using “tangents” $(\bm{v},1)$
A.2 Norm L2 Loss
@@
$L_{\text{norm}}(\bm{p},\bm{y})=\frac{e}{\sqrt{e+c}}$ (18)
where $c$ is a small constant. This formulation is equivalent to the adaptive L2 loss proposed in MeanFlow (Geng et al., 2025) with $p=0.5$, and behaves similarly to a Huber loss, being robust to large errors.
A.3 Experimental Details
@@
(b) Sampling Strategy. Our multi-step sampling (NFE $\geq$ 2) follows a standard iterative refinement process. For an $N$-step generation, we use a simple schedule of $N$ equally spaced timesteps $t_{i}=(i-1)/N$ for $i=1,\dots,N$. The process starts with pure noise $\bm{x}_{0}$. At each step $i$, we first compute a 1-step prediction $\hat{\bm{x}}_{1}$ using the model’s output $\bm{F}_{\theta}$: $\hat{\bm{x}}_{1}=\bm{x}_{t_{i}}+(1-t_{i})\bm{F}_{\theta}(\bm{x}_{t_{i}},c_{\text{CM}})$. If it is not the final step, we generate the input for the next step, $\bm{x}_{t_{i+1}}$, by linearly interpolating between the predicted endpoint and a new noise sample, consistent with the OT-FM framework:
@@
(c) Reproduction Details. As the official code for MeanFlow (Geng et al., 2025) and sCM (Lu & Song, 2024) was not available, the results in our tables are from our own reproductions of their papers. Our MeanFlow implementation follows its two-time-variable conditioning and log-normal time sampling. Following their from-scratch regime, we set $t=r$ with a 75% probability for optimal performance. In the distillation paradigm, we found this setting struggled to converge and therefore did not use it. For sCM, we incorporated all necessary techniques described in their work, including pixel normalization, tangent warmup, tangent normalization, and adaptive weighting, to ensure stable training. We did not use the “TrigFlow” proposed in sCM, as we believe the specific flow construction is not key to building continuous consistency models. To ensure reproducibility, our open-source code will include our implementations of these baselines.
(e) Time Sampling Schedule. Following sCM (Lu & Song, 2024), the time $t\in[0,1]$ is sampled according to a schedule that concentrates samples near the data endpoint ($t=1$). We first sample a value $\sigma$ from a log-normal distribution, i.e., $\ln(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$, and then compute $t$ as:
@@
A.4 Hyperparameters

### v2

@@ 5 Limitations and Future Work
Our work highlights two primary areas for future research. First, on large-scale models, a performance gap persists between samples generated in minimal steps (e.g., 1-2) and those requiring more steps (e.g., 8). Bridging this gap by enhancing the model’s expressiveness in the ultra-few-step regime is a key challenge. Second, while our Chain-JVP method successfully mitigates the memory bottleneck of the Jacobian-vector product, its computational overhead remains a concern. Optimizing its efficiency is crucial for improving training throughput. Additionally, we found that our acceleration model, even when fine-tuned exclusively on T2I data, can directly accelerate T2V synthesis by targeting only the low-noise diffusion steps ($\text{SNR}\leq\frac{\text{SNR}_{\text{min}}}{2}$), all without introducing flickering or detail loss. This could inform future work on efficient, high-fidelity video synthesis.
@@ 6 Conclusion
[…]
@@ Appendix A Appendix
A.1 Theoretical Analysis: Stability and Convergence of FACM
This section provides a detailed mathematical derivation of the stability and convergence properties of FACM, complementing the intuitive discussion in Section 3.
We use the same notation as in the main text:
  • Student network (online model): $\bm{F}_{\theta}(\bm{x}_{t},t)$
  • Stop-gradient copy: $\bm{F}_{\theta^{-}}(\bm{x}_{t},t)=\text{sg}(\bm{F}_{\theta}(\bm{x}_{t},t))$
  • Instantaneous velocity field (FM anchor): $\bm{v}(\bm{x}_{t},t)$
  • CM operator (Consistency target): $\mathcal{T}[\bm{F}]=\bm{v}+(1-t)\frac{\mathrm{d}\bm{F}}{\mathrm{d}t}$
A.1.1 Stability Analysis: Prevention of Gradient Explosion
Step 1.1: Structural form of the CM gradient.
The CM loss is
$\mathcal{L}_{\text{CM}}(\theta)=\mathbb{E}_{\bm{x}_{t},t}\left[\frac{1}{2}\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\big\|^{2}\right],$ (16)
where $\bm{v}_{\text{tar}}$ is the CM target derived from the operator $\mathcal{T}$ (Eq. 13 and Eq. 2 in the main text).
Differentiating w.r.t. $\theta$ yields the gradient in compact form:
$\nabla_{\theta}\mathcal{L}_{\text{CM}}$ $=\nabla_{\theta}\mathbb{E}_{\bm{x}_{t},t}\left[\frac{1}{2}\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\big\|^{2}\right]$ (17)
$=\mathbb{E}_{\bm{x}_{t},t}\left[\nabla_{\theta}\left(\frac{1}{2}(\bm{F}_{\theta}-\bm{v}_{\text{tar}})^{\top}(\bm{F}_{\theta}-\bm{v}_{\text{tar}})\right)\right]$
$=\mathbb{E}_{\bm{x}_{t},t}\left[\left(\nabla_{\theta}\bm{F}_{\theta}-\nabla_{\theta}\bm{v}_{\text{tar}}\right)^{\top}(\bm{F}_{\theta}-\bm{v}_{\text{tar}})\right]$
(Since $\bm{v}_{\text{tar}}$ is a stop-gradient target, $\nabla_{\theta}\bm{v}_{\text{tar}}=0$)
$=\mathbb{E}_{\bm{x}_{t},t}\Big[\underbrace{\nabla_{\theta}\bm{F}_{\theta}(\bm{x}_{t},t)^{\top}}_{\text{parameter sensitivity}}\cdot\underbrace{\big(\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\big)}_{\text{prediction error }\bm{e}}\Big],$
where $\nabla_{\theta}\bm{F}_{\theta}$ denotes the Jacobian of $\bm{F}_{\theta}$ w.r.t. the parameters $\theta$, and we write
$\bm{e}(\bm{x}_{t},t;\theta):=\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t).$ (18)
Taking norms and using $\|A^{\top}b\|\leq\|A\|_{\text{op}}\,\|b\|$, we obtain
$\big\|\nabla_{\theta}\mathcal{L}_{\text{CM}}\big\|\leq\mathbb{E}_{\bm{x}_{t},t}\Big[\big\|\nabla_{\theta}\bm{F}_{\theta}(\bm{x}_{t},t)\big\|_{\text{op}}\cdot\|\bm{e}(\bm{x}_{t},t;\theta)\|\Big].$ (19)
Thus, the CM gradient norm is governed by two independent factors:
  • the prediction error $\bm{e}$, which determines the basic scale and direction of the gradient;
  • the parameter sensitivity $\nabla_{\theta}\bm{F}_{\theta}$, which acts as a multiplicative amplifier.
Step 1.2: Decomposition of the error term.
Recall that the CM operator for a general field $\bm{F}$ is
$\bm{v}_{\text{tar}}(\bm{x}_{t},t)=\bm{v}(\bm{x}_{t},t)+(1-t)\Big(\partial_{t}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)+\nabla_{\bm{x}_{t}}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)\cdot\bm{v}(\bm{x}_{t},t)\Big).$ (20)
The error of the online model $\bm{F}_{\theta}$ relative to this target then decomposes as
$\bm{e}=\underbrace{\big(\bm{F}_{\theta}-\bm{v}\big)}_{\text{function deviation}}\;-\;\underbrace{(1-t)\frac{\partial\bm{F}_{\theta^{-}}}{\partial t}}_{\text{time derivative term}}\;-\;\underbrace{(1-t)\nabla_{\bm{x}_{t}}\bm{F}_{\theta^{-}}\cdot\bm{v}}_{\text{JVP (spatial Jacobian)}}.$ (21)
Hence, the size of $\bm{e}$ is governed by the first-order spatio-temporal derivatives of the (stop-gradient) network $\bm{F}_{\theta^{-}}$.
Step 1.3: FACM’s stabilization mechanism.
FACM combines Flow Matching (FM) with the CM objective, using shared parameters $\theta$, and thereby stabilizes both factors in Eq. (19).
(1) Lipschitz supervision via Flow Matching. The FM loss (Eq. 9) trains $\bm{F}_{\theta}(\bm{x}_{t},c_{\text{FM}})$ to match the instantaneous velocity field $\bm{v}(\bm{x}_{t},t)$, which is a bounded ground-truth function that does not depend on $\theta$ and is Lipschitz in $(\bm{x}_{t},t)$. For standard architectures, the Lipschitz constant of $\bm{F}_{\theta}$ with respect to its inputs is determined only by the spectral norms of the weight matrices and the activation Lipschitz constants, all of which are shared across time conditions $t$ and 2t2-t. Minimizing the FM loss therefore keeps these spectral norms in a moderate range and induces a global Lipschitz bound
$\big\|\nabla_{\theta}\bm{F}_{\theta}(\bm{x}_{t},c)\big\|_{\text{op}}\leq L_{\text{net}}$ (22)
for all $(\bm{x}_{t},c)$ in the training domain, including both the FM branch (c=2tc=2-t) and the CM branch ($c=t$). Here, $L_{\text{net}}$ represents the Lipschitz constant of the network, and $\|\cdot\|_{\text{op}}$ denotes the spectral norm (operator norm). In contrast, pure CM supervises $\bm{F}_{\theta}$ with a self-referential target. To satisfy the consistency boundary condition (i.e., mapping $\bm{x}_{t}$ to $\bm{x}_{1}$ as $t\to 1$) under the parameterization $\bm{x}_{t}+(1-t)\bm{F}_{\theta}(\bm{x}_{t},t)$, the network output $\bm{F}_{\theta}$ is implicitly forced to approximate the average velocity $(\bm{x}_{1}-\bm{x}_{t})/(1-t)$, which blows up as $t\to 1$. The CM target is therefore a dynamic, potentially unbounded prediction, whereas FM always provides a bounded ground-truth target.
(2) Bounding the error via FM-anchored supervision. Next, we show that the prediction error term $\bm{e}$ in the gradient (Eq. 19) remains bounded under FACM. For clarity, we focus on the deviation between the online model and the FM anchor, $\bm{F}_{\theta}(\bm{x}_{t},t)$ vs. $\bm{v}(\bm{x}_{t},t)$, and make explicit use of the expanded time interval strategy (Figure 2), where FM uses the condition $c_{\text{FM}}=2-t$ but predicts the same physical velocity $\bm{v}(\bm{x}_{t},t)$.
We begin with the triangle inequality:
$\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}(\bm{x}_{t},t)\big\|$ $\leq\underbrace{\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta}(\bm{x}_{t},2-t)\big\|}_{\text{temporal smoothness of }\bm{F}_{\theta}}$ (23)
$\quad+\underbrace{\big\|\bm{F}_{\theta}(\bm{x}_{t},2-t)-\bm{v}(\bm{x}_{t},t)\big\|}_{\text{FM error}}.$
The first term measures how much the network output changes when the (time-related) condition goes from $t$ to 2t2-t for the same spatial point $\bm{x}_{t}$. Using the fundamental theorem of calculus for the time coordinate, we write
$\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta}(\bm{x}_{t},2-t)=\int_{2-t}^{t}\frac{\partial}{\partial\tau}\bm{F}_{\theta}(\bm{x}_{t},\tau,\tilde{c}(\tau))\,d\tau,$ (24)
where $\tilde{c}(\tau)$ interpolates between the CM and FM conditions as $\tau$ varies. Taking norms and applying the triangle inequality gives
$\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta}(\bm{x}_{t},2-t)\big\|\leq\left|\int_{2-t}^{t}\Big\|\frac{\partial}{\partial\tau}\bm{F}_{\theta}(\bm{x}_{t},\tau,\tilde{c}(\tau))\Big\|d\tau\right|.$ (25)
Since FM constrains the spectral norms of the weights, the partial time derivative $\big\|\partial_{\tau}\bm{F}_{\theta}(\cdot)\big\|$ is bounded by a constant (of order $L_{\text{net}}$) over the training domain. The integration interval has length at most $2$, so
$\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta}(\bm{x}_{t},2-t)\big\|\leq 2L_{\text{net}}.$ (26)
The second term in Eq. (23) is precisely the FM error
$\varepsilon_{\text{FM}}(\bm{x}_{t},t):=\bm{F}_{\theta}(\bm{x}_{t},2-t)-\bm{v}(\bm{x}_{t},t),$ (27)
whose squared norm is minimized by $\mathcal{L}_{\text{FM}}$ and thus has bounded variance. Combining these bounds, we obtain a uniform control of the function deviation:
$\big\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}(\bm{x}_{t},t)\big\|\leq 2L_{\text{net}}+\|\varepsilon_{\text{FM}}(\bm{x}_{t},t)\|.$ (28)
Finally, because the same spectral-norm constraints apply to all weight matrices, the spatial Jacobian $\nabla_{\bm{x}_{t}}\bm{F}_{\theta^{-}}$ and time derivative $\partial_{t}\bm{F}_{\theta^{-}}$ in Eq. (21) are also uniformly bounded by constants of order $L_{\text{net}}$. Hence, all components of $\bm{e}$ remain bounded.
Conclusion (Stability).
Putting these results together, we see that under FACM:
  • the parameter sensitivity is bounded: $\|\nabla_{\theta}\bm{F}_{\theta}\|_{\text{op}}\leq L_{\text{net}}$;
  • the prediction error $\bm{e}$ is uniformly bounded in norm by a constant depending on $L_{\text{net}}$ and the FM error statistics.
Therefore, the CM gradient norm satisfies
$\big\|\nabla_{\theta}\mathcal{L}_{\text{CM}}\big\|\leq C\,L_{\text{net}}^{2},$ (29)
eliminating the gradient explosion observed in pure CM training.
A.1.2 Convergence Proof
We provide a concise analysis showing that FACM ensures convergence by eliminating target singularities, bounding gradient variance, and enforcing alignment through shared parameters.
Step 2.1: Mechanism of Variance Reduction.
The pure consistency target implies predicting the average velocity
$\overline{\bm{v}}(\bm{x}_{t},t)=\frac{\bm{x}_{1}-\bm{x}_{t}}{1-t}.$ (30)
As $t\to 1$, any variance $\sigma^{2}$ in the endpoint estimate $\bm{x}_{1}$ is amplified by $(1-t)^{-2}$, so
$\mathrm{Var}[\overline{\bm{v}}(\bm{x}_{t},t)\mid t]=\frac{\sigma^{2}}{(1-t)^{2}},$ (31)
which makes the pure-CM shortcut objective ill conditioned near the data endpoint. FACM mitigates this by using a relaxed target
$\bm{v}_{\text{tar}}(\bm{x}_{t},t)=(1-\alpha(t))\bm{F}_{\theta^{-}}(\bm{x}_{t},t)+\alpha(t)\,\overline{\bm{v}}(\bm{x}_{t},t).$ (32)
In practice we use the schedule $\alpha(t)=1-t^{0.5}$ (Sec. 3). Writing $t=1-\varepsilon$ with $0<\varepsilon\ll 1$ and using the Taylor expansion
$t^{0.5}=(1-\varepsilon)^{0.5}=1-\tfrac{1}{2}\varepsilon+\mathcal{O}(\varepsilon^{2}),$ (33)
we obtain
$\alpha(t)=1-t^{0.5}=\tfrac{1}{2}\varepsilon+\mathcal{O}(\varepsilon^{2})=\tfrac{1}{2}(1-t)+\mathcal{O}\big((1-t)^{2}\big).$ (34)
Hence $\alpha(t)$ is asymptotically proportional to (1t)(1-t) and
$\lim_{t\to 1}\frac{\alpha(t)}{1-t}=\tfrac{1}{2},$ (35)
so the factor $\alpha(t)$ cancels the $(1-t)^{-1}$ singularity in the average-velocity term up to a constant. Consequently $\mathrm{Var}[\bm{v}_{\text{tar}}(\bm{x}_{t},t)\mid t]$ remains uniformly bounded over $t\in[0,1]$, providing the core mechanism for variance reduction in FACM.
Step 2.2: Bounded and Reduced Gradient Variance.
Locally, $L_{\text{norm}}$ behaves like a rescaled squared $\ell_{2}$ loss, so the CM gradient for one sample $(\bm{x}_{t},t)$ satisfies
$\big\|\nabla_{\theta}\ell_{\text{CM}}(\theta;\bm{x}_{t},t)\big\|\lesssim\beta(t)\,\|\nabla_{\theta}\bm{F}_{\theta}(\bm{x}_{t},t)\|_{\text{op}}\,\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\|^{2}.$ (36)
Using the stability bound $\|\nabla_{\theta}\bm{F}_{\theta}\|_{\text{op}}\leq L_{\text{net}}$ and the uniform boundedness of $\bm{v}_{\text{tar}}$, we obtain a finite second-moment (and hence variance) bound
$\mathbb{E}\big[\|\widehat{\nabla}_{\theta}\mathcal{L}_{\text{CM}}\|^{2}\big]\;\lesssim\;L_{\text{net}}^{2}\,\mathbb{E}_{t}\Big[\beta(t)^{2}\,\mathbb{E}_{\bm{x}_{t}}\big[\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\|^{2}\mid t\big]\Big].$ (37)
Here the boundedness follows from the variance-reduction mechanism in the previous paragraph, which shows that $\mathrm{Var}[\bm{v}_{\text{tar}}(\bm{x}_{t},t)\mid t]$ is uniformly bounded in $t$, together with the global Lipschitz bound $\|\nabla_{\theta}\bm{F}_{\theta}\|_{\text{op}}\leq L_{\text{net}}$ from Step 1.4: even if $\beta(t)\equiv 1$, the right-hand side is finite because both the Jacobian norm and the target variance are controlled. The role of $\beta(t)\in[0,1]$ is to further reduce the integrated variance: for any non-negative function $q(t)$,
$\mathbb{E}_{t}[\beta(t)^{2}q(t)]\leq\mathbb{E}_{t}[q(t)],$ (38)
with strict inequality whenever $\beta(t)<1$ on a set of non-zero measure. Since $q(t)=\mathbb{E}_{\bm{x}_{t}}\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{v}_{\text{tar}}(\bm{x}_{t},t)\|^{2}\mid t$ is typically largest near the data endpoint $t\approx 1$, choosing a decaying schedule (e.g., cosine) for $\beta(t)$ suppresses precisely those high-variance contributions, yielding a strictly lower integrated gradient variance than the pure-CM case.
Step 2.3: Alignment via Shared Parameters.
The expanded time interval $[0,2]$ creates a natural synchronization mechanism between the FM and CM tasks. In the high-SNR region ($t\approx 1$), the schedules $\alpha(t)$ and $\beta(t)$ suppress the CM gradients, so parameter updates there are dominated by the FM branch, which supervises $\bm{F}_{\theta}(\bm{x}_{t},2-t)$ to match the instantaneous velocity field $\bm{v}(\bm{x}_{t},t)$. Because both branches share the same parameters and the stability analysis above bounds the discrepancy between $\bm{F}_{\theta}(\bm{x}_{t},t)$ and $\bm{F}_{\theta}(\bm{x}_{t},2-t)$, this FM supervision implicitly guides $\bm{F}_{\theta}(\bm{x}_{t},t)$ to align with the underlying velocity field. Combined with the reduced gradient variance from Step 2.2, this shared-parameter coupling explains the empirically faster and more stable convergence of FACM compared to pure CM training.
A.2 On Total Derivatives
@@
In this paper, for a network $N(\bm{x}_{t},\mathcal{C}(t))$ (e.g., $\bm{F}_{\theta}$), its total derivative along the trajectory $\bm{x}_{t}(t)=(1-t)\bm{x}_{0}+t\bm{x}_{1}$ (with $\bm{v}=\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}=\bm{x}_{1}-\bm{x}_{0}$) with respect to $t$ is given by the chain rule:
$\frac{\mathrm{d}N(\bm{x}_{t},\mathcal{C}(t))}{\mathrm{d}t}=\frac{\partial N}{\partial\bm{x}_{t}}\bm{v}+\nabla_{\mathcal{C}}N\cdot\frac{\mathrm{d}\mathcal{C}(t)}{\mathrm{d}t}.$ (39)
@@
The term $\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}$ is computed for the CM task. Depending on the implementation strategy (Sec. 3.3.2), the conditioning $c_{\text{CM}}$ can be $t$ or a tuple $(t,1)$. In both cases, its derivative with respect to $t$ is effectively 1 for the time-dependent component and 0 for any constant component. Therefore, the calculation simplifies to:
$\frac{\mathrm{d}N(\bm{x}_{t},c_{\text{CM}})}{\mathrm{d}t}\approx\frac{\partial N}{\partial\bm{x}_{t}}\bm{v}+\frac{\partial N}{\partial t},$ (40)
@@
where $\frac{\partial N}{\partial t}$ denotes the partial derivative with respect to the explicit time argument(s) encoded in the conditioning.
A.3 Norm L2 Loss
@@
$L_{\text{norm}}(\bm{p},\bm{y})=\frac{e}{\sqrt{e+c^{2}}}$ (41)
where $c$ is a small constant. This formulation is equivalent to the adaptive L2 loss proposed in MeanFlow (geng2025meanflowsonestepgenerative) with $p=0.5$, and behaves similarly to a Huber loss, being robust to large errors.
A.4 Experimental Details
@@
(b) Sampling Strategy. Our multi-step sampling (NFE $\geq$ 2) follows a standard iterative refinement process. For an $N$-step generation, we use a simple schedule of $N$ equally spaced timesteps $t_{i}=(i-1)/N$ for $i=1,\dots,N$. The process starts with pure noise $\bm{x}_{0}$. At each step $i$, we first compute a one-step prediction $\hat{\bm{x}}_{1}$ using the model’s output $\bm{F}_{\theta}$: $\hat{\bm{x}}_{1}=\bm{x}_{t_{i}}+(1-t_{i})\bm{F}_{\theta}(\bm{x}_{t_{i}},c_{\text{CM}})$. If it is not the final step, we generate the input for the next step, $\bm{x}_{t_{i+1}}$, by linearly interpolating between the predicted endpoint and a new noise sample, consistent with the OT-FM framework:
@@
(c) Reproduction Details. At the time of our main ablations, the official codebases for MeanFlow (geng2025meanflowsonestepgenerative) and sCM (lu2024simplifying) were not yet available. A JAX implementation of MeanFlow was later released, but without a reproducible configuration for its SOTA results. For a controlled and fair comparison, we therefore implemented PyTorch reproductions under the exact same environment, teacher, and hyperparameters across methods. Our MeanFlow reproduction follows its two-time-variable conditioning and log-normal time sampling; following the from-scratch regime, we set $t=r$ with a 75% probability for optimal performance. In the distillation setting, this configuration struggled to converge and was therefore not used. For sCM, we incorporated all necessary techniques described in their work, including pixel normalization, tangent warmup, tangent normalization, and adaptive weighting, to ensure stable training. We did not use the TrigFlow proposed in sCM, as we believe the specific flow construction is orthogonal to building continuous-time consistency models. We will release our reproductions alongside our code to ensure full reproducibility.
(d) Classifier-Free Guidance in Distillation. When distilling a teacher model that supports classifier-free guidance (CFG), we compute both the conditional velocity $\bm{v}_{\text{cond}}$ and unconditional velocity $\bm{v}_{\text{uncond}}$ from the teacher, and construct the target as
$\bm{v}=\bm{v}_{\text{uncond}}+w\cdot(\bm{v}_{\text{cond}}-\bm{v}_{\text{uncond}}),$ (43)
where $w$ is the guidance weight. During training, the unconditional forward is computed with torch.no_grad(), adding less than 5% overhead, which is consistent with sCM and MeanFlow. To stabilize the high-noise region, we set a time threshold $t_{\text{low}}$ and disable guidance for $t<t_{\text{low}}$ (we use $t_{\text{low}}{=}0.05$ in our experiments). During pre-training, the null-token probability is 10%, and the condition is not dropped during distillation. At inference, FACM uses a single timestamp; even under CFG, each step requires only one NFE.
(e) Time Sampling Schedule. Following sCM (lu2024simplifying), the time $t\in[0,1]$ is sampled according to a schedule that concentrates samples near the data endpoint ($t=1$). We first sample a value $\sigma$ from a log-normal distribution, i.e., $\ln(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$, and then compute $t$ as:
@@
(f) Weighting Functions. For the CM loss component (Eq. 13), we find that the weighting functions $\alpha(t)=1-t^{0.5}$ and $\beta(t)=\cos(t\cdot\pi/2)$ provide an effective general solution. These functions are crucial for navigating the trade-off between ensuring endpoint quality (in high-SNR regions) and satisfying global consistency (in low-SNR regions).
A.5 Discussion: From-Scratch Training vs. Distillation
While our method can be trained from scratch and achieves a competitive result (See Table 3), we identify the two-stage distillation paradigm as the more principled and practically superior approach. Attempting to learn both the anchor and the shortcut simultaneously from scratch introduces a “chicken-and-egg” problem, as the model must learn a shortcut based on a trajectory it has not yet accurately modeled. This creates an unstable “moving target” for optimization and incurs higher computational costs. In contrast, distillation from a pre-trained FM teacher provides a fixed, high-quality velocity field, offering a much more stable and well-defined learning objective. MeanFlow (geng2025meanflowsonestepgenerative) also encounters this problem in the from-scratch setting, achieving its optimal performance only by having its objective degenerate to a Flow Matching task for a large portion of samples (e.g., 75%), which further validates our core thesis that a robust foundation in the velocity field is a prerequisite for learning stable shortcuts.
A.6 Flow Mapping Equivalence Derivation
Let the Flow Mapping function be $f_{\theta}(\bm{x}_{t},t,r)=\bm{x}_{t}+(r-t)\bm{F}_{\theta}(\bm{x}_{t},t,r)$. The consistency condition $\frac{d}{dt}f_{\theta}(\bm{x}_{t},t,r)=0$ is equivalent to the learning objective for $\bm{F}_{\theta}$:
$\frac{d}{dt}f_{\theta}(\bm{x}_{t},t,r)=0\quad$ $\iff\quad\bm{v}-\bm{F}_{\theta}(\bm{x}_{t},t,r)+(r-t)\frac{\mathrm{d}\bm{F}_{\theta}}{\mathrm{d}t}=0$
$\iff\quad\bm{F}_{\theta}(\bm{x}_{t},t,r)=\bm{v}+(r-t)\frac{\mathrm{d}\bm{F}_{\theta}}{\mathrm{d}t}$
Enforcing this for $t\in[0,r]$ implies that $f_{\theta}$ is constant over the interval, thus mapping any point $\bm{x}_{t}$ to the endpoint $\mathbf{x}_{r}$:
$f_{\theta}(\bm{x}_{t},t,r)\stackrel{{\scriptstyle t\in[0,r]}}{{=}}f_{\theta}(\mathbf{x}_{r},r,r)=\mathbf{x}_{r}$
A.7 Hyperparameters